# Mirror and Preconditioned Gradient Descent in Wasserstein Space

### v1

@@ C.1 Wasserstein differentials
We recall the notion of Wasserstein differentiability introduced in [17, 69]. First, we introduce sub and super differential.
@@
Let $\mathcal{F}:\mathcal{P}_{2}(\mathbb{R}^{d})\to(-\infty,+\infty]$ lower semi-continuous and denote $D(\mathcal{F})=\{\mu\in\mathcal{P}_{2}(\mathbb{R}^{d}),\ \mathcal{F}(\mu)<\infty\}$. Let $\mu\in D(\mathcal{F})$. Then, a map $\xi\in L^{2}(\mu)$ belongs to the subdifferential $\partial^{-}\mathcal{F}(\mu)$ of $\mathcal{F}$ at $\mu$ if for all $\nu\in\mathcal{P}_{2}(\mathbb{R}^{d})$,
@@
$\mathcal{T}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d})=\overline{\left\{\nabla\psi,\;\psi\in\mathcal{C}_{c}^{\infty}(\mathbb{R}^{d})\right\}}\subset L^{2}(\mu),$
where the closure is taken in $L^{2}(\mu)$, see Ambrosio et al. [4, Definition 8.4.1]. Lanzetti et al. [69, Proposition 2.5] showed that if $\mathcal{F}$ is Wasserstein differentiable, then there is always a unique gradient living in the tangent space, and we can restrict ourselves without loss of generality to this gradient.
@@
The Wasserstein gradient of $\mathcal{F}$ can be computed in practice using the first variation $\frac{\delta\mathcal{F}}{\delta\mu}$ [102, Definition 7.12], which is defined, if is exists, as the unique function (up to a constant) such that, for $\chi$ satisfying $\int\mathrm{d}\chi=0$,
@@
Then the Wasserstein gradient can be computed as $\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mu)=\nabla\frac{\delta\mathcal{F}}{\delta\mu}(\mu)$.
@@
$\quad+o\left(\sqrt{\int\|x-y\|_{2}^{2}\ \mathrm{d}\gamma(x,y)}\right)$
Thus, $\delta\tilde{\mathcal{F}}_{\mu}(\mathrm{S},\mathrm{T}-\mathrm{S})=\langle\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mathrm{S}_{\#}\mu)\circ\mathrm{S},\mathrm{T}-\mathrm{S}\rangle_{L^{2}(\mu)}$. Note that in the third equality we used that $\nabla_{\mathrm{W}_{2}}\mathcal{F}\in L^{2}(\mu)$. ∎
A similar formula can be found in Gangbo and Tudorascu [51, Corollary 3.22], however the space $H$ used there is not $L^{2}(\mu)$ but a lifting $L^{2}(\Omega;\mathbb{R}^{d})$ of measures on random variables. They should not be confused.
@@ C.2 Wasserstein Hessians
A natural object of interest is the Hessian of the objective $\mathcal{F}$, which we define below. This notion is usually defined along Wasserstein geodesics, i.e. curves of the form $\mu_{t}=(\mathrm{Id}+t\nabla\psi)_{\#}\mu$ for $\psi\in\mathcal{C}_{c}^{\infty}(\mathbb{R}^{d})$ [67].
@@
The Wasserstein Hessian of a functional $\mathcal{F}:\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}$ at $\mu$ is defined for any $\psi\in\mathcal{C}_{c}^{\infty}(\mathbb{R}^{d})$ as:
[…]
where $(\mu_{t},v_{t})_{t\in[0,1]}$ is a Wasserstein geodesic with $\mu_{0}=0,v_{0}=\nabla\psi$.
Definition 8 can be straightforwardly related to the usual symmetric bilinear form defined on $\mathcal{T}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d})\times\mathcal{T}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d})$ [90, Section 3]:
Definition 9.
In this work, we are interested in more general curves, which are acceleration free, i.e. of the form $\mu_{t}=(\mathrm{S}+tv)_{\#}\mu$ with $\mathrm{S},v\in L^{2}(\mu)$. Thus, we define analogously the Hessian and the Hessian operator along such curves. We note that if $\mathrm{S}$ is invertible, $\mu_{t}=(\mathrm{Id}+tv\circ\mathrm{S}^{-1})_{\#}\mathrm{S}_{\#}\mu$, and the notions can be linked with Wasserstein Hessian. However, in general, this does not need to be the case.
Definition 10.
The Hessian of $\mathcal{F}:\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}$ along $t\mapsto\mu_{t}=(\mathrm{S}+tv)_{\#}\mu$ for $\mathrm{S},v\in L^{2}(\mu)$ is defined as
$\operatorname{Hess}_{\mu,t}\mathcal{F}(v,v)=\frac{\mathrm{d}^{2}}{\mathrm{d}t^{2}}\mathcal{F}(\mu_{t}).$ (28)
Moreover, we define the Hessian operator $\mathrm{H}\mathcal{F}_{\mu,t}:L^{2}(\mu)\to L^{2}(\mu)$ as the operator satisfying for all $t\in[0,1]$,
$\frac{\mathrm{d}^{2}}{\mathrm{d}t^{2}}\mathcal{F}(\mu_{t})=\langle\mathrm{H}\mathcal{F}_{\mu,t}v,v\rangle_{L^{2}(\mu)}.$ (29)
Wang and Li [115] derived a general closed form of the Wasserstein Hessian through the first variation of $\mathcal{F}$. Here, we extend their formula along any curve $\mu_{t}=(\mathrm{S}+tv)_{\#}\mu$ with $\mathrm{S},v\in L^{2}(\mu)$. We first provide a lemma computing the derivative of the Wasserstein gradient.
@@
Let $\mathcal{F}:\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}$ be twice continuously differentiable and assume that $\frac{\delta}{\delta\mu}\nabla\frac{\delta\mathcal{F}}{\delta\mu}=\nabla\frac{\delta^{2}\mathcal{F}}{\delta\mu^{2}}$. Let $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$ and for all $t\in[0,1]$, $\mu_{t}=(\mathrm{T}_{t})_{\#}\mu$ where $\mathrm{T}_{t}$ is differentiable w.r.t. $t$ with $\frac{\mathrm{d}\mathrm{T}_{t}}{\mathrm{d}t}\in L^{2}(\mu)$. Then,
@@
Under the same assumptions as in Lemma 9, let $\mu_{t}=(\mathrm{T}_{t})_{\#}\mu$ with $\mathrm{T}_{t}=\mathrm{S}+tv$, $\mathrm{S},v\in L^{2}(\mu)$, then
with $\mathrm{H}\mathcal{F}_{\mu,t}:L^{2}(\mu)\to L^{2}(\mu)$ defined as, for all $v\in L^{2}(\mu)$, $x\in\mathbb{R}^{d}$,
@@
We note that if $\mathrm{T}_{t}$ is invertible for all $t$, with $v_{t}=v\circ\mathrm{T}_{t}^{-1}$, we can write
@@
$=\int\mathrm{H}\mathcal{F}_{\mu,t}[v](x),v(x)\rangle\ \mathrm{d}\mu(x)$
@@
$=\langle\mathrm{H}\mathcal{F}_{\mu_{t}}v_{t},v_{t}\rangle_{L^{2}(\mu_{t})},$
because
$\mathrm{H}\mathcal{F}_{\mu,t}[v](x)$ $=\int\nabla_{y}\nabla_{x}\frac{\delta^{2}\mathcal{F}}{\delta\mu^{2}}(\mu_{t})(\mathrm{T}_{t}(x),y_{t})v_{t}(y_{t})\ \mathrm{d}\mu_{t}(y)+\nabla^{2}\frac{\delta\mathcal{F}}{\delta\mu}(\mu_{t})\big{(}\mathrm{T}_{t}(x)\big{)}v(x),$ (34)
and thus
$\mathrm{H}\mathcal{F}_{\mu,t}[v]\big{(}\mathrm{T}_{t}^{-1}(x_{t})\big{)}$ $=\int\nabla_{y}\nabla_{x}\frac{\delta^{2}\mathcal{F}}{\delta\mu^{2}}(\mu_{t})(x_{t},y_{t})v_{t}(y_{t})\ \mathrm{d}\mu_{t}(y)+\nabla^{2}\frac{\delta\mathcal{F}}{\delta\mu}(\mu_{t})(x_{t})v_{t}(x_{t})$ (35)
$=\mathrm{H}\mathcal{F}_{\mu_{t}}[v_{t}](x_{t}).$
@@
Let $\mathcal{V}(\mu)=\int V\ \mathrm{d}\mu$ with $V$ convex and twice differentiable. Then, it is well known that $\frac{\delta\mathcal{V}}{\delta\mu}(\mu)=V$ and $\frac{\delta^{2}\mathcal{V}}{\delta\mu^{2}}=0$. Thus, applying Proposition 10, we recover for $\mu_{t}=(\mathrm{T}_{t})_{\#}\mu$,
@@
Let $\mathcal{W}(\mu)=\iint W(x-y)\ \mathrm{d}\mu(x)\mathrm{d}\mu(y)$ with $W$ convex, symmetric and twice differentiable. Then, we have for all $x,y\in\mathbb{R}^{d}$, $\frac{\delta\mathrm{W}}{\delta\mu}(x)=(W\star\mu)(x)$ and $\frac{\delta^{2}\mathrm{W}}{\delta\mu^{2}}(x,y)=W(x-y)$ (see e.g. [115, Example 7]), and thus applying Proposition 10, for $\mu_{t}=(\mathrm{T}_{t})_{\#}\mu$, the operator is
@@ C.3 Convexity in Wasserstein space
We first recall the definition of $\alpha$-convex functionals [4, Definition 9.1.1].
@@
Let $\mathcal{F}:\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}$ be twice continuously differentiable. Let $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$, $\mathrm{T},\mathrm{S}\in L^{2}(\mu)$, $\mu_{t}=\big{(}\mathrm{T}_{t}\big{)}_{\#}\mu$ for all $t\in[0,1]$ where $\mathrm{T}_{t}=(1-t)\mathrm{S}+t\mathrm{T}$. Furthermore, denote for $t_{1},t_{2}\in[0,1]$, $\tilde{\mu}_{t}^{t_{1}\to t_{2}}=\big{(}(1-t)\mathrm{T}_{t_{1}}+t\mathrm{T}_{t_{2}}\big{)}_{\#}\mu$. Then, the following statement are equivalent,
@@
If we require the convexity and the smoothness to hold along any curve of the form $\mu_{t}=\big{(}(1-t)\mathrm{S}+t\mathrm{T})_{\#}\mu$ for $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$ and $\mathrm{T},\mathrm{S}\in L^{2}(\mu)$, then it coincides with the transport convexity and smoothness recently introduced by Tanaka [109, Definitions 4.1 and 4.5] as by Proposition 8, $\delta\tilde{\mathcal{F}}_{\mu}(\mathrm{S},\mathrm{T}-\mathrm{S})=\langle\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mathrm{S}_{\#}\mu)\circ\mathrm{S},\mathrm{T}-\mathrm{S}\rangle_{L^{2}(\mu)}$, and thus the convexity reads as follows
@@
This type of convexity is actually a particular case of the notion of convexity along acceleration-free curves introduced by Parker [91] (also introduced by Cavagnari et al. [27] under the name of total convexity). The latter requires convexity to hold along any curve of the form $\mu_{t}=\big{(}(1-t)\pi^{1}+t\pi^{2}\big{)}_{\#}\gamma$ with $\gamma\in\Pi(\mu,\nu)$, $\mu,\nu\in\mathcal{P}_{2}(\mathbb{R}^{d})$ and $\pi^{1}(x,y)=x$, $\pi^{2}(x,y)=y$. The transport convexity of Tanaka [109] is thus a particular case for couplings obtained through maps, i.e. $\gamma=(\mathrm{T},\mathrm{S})_{\#}\mu$. Parker [91] notably showed that this notion of convexity is equivalent with the geodesic convexity for Wasserstein differentiable functionals.
We can also define the strict convexity using strict inequalities in Proposition 11-(c1)-(c2)-(c3) (but not in (c4)).
Finally, as we defined the relative $\alpha$-convexity and $\beta$-smoothness of $\mathcal{F}$ relative to $\mathcal{G}$ using Bregman divergences in Definition 3, we can show that it is equivalent with $\mathcal{F}-\alpha\mathcal{G}$ and $\beta\mathcal{G}-\mathcal{F}$ being convex.
@@
By Definition 3, $\mathcal{F}$ is $\alpha$-convex relative to $\mathcal{G}$ along $t\mapsto\mu_{t}$ if for all $s,t\in[0,1]$, $\mathrm{d}_{\tilde{\mathcal{F}}_{\mu}}(\mathrm{T}_{s},\mathrm{T}_{t})\geq\alpha\mathrm{d}_{\tilde{\mathcal{G}}_{\mu}}(\mathrm{T}_{s},\mathrm{T}_{t})$. This is equivalent with $\mathrm{d}_{\tilde{\mathcal{F}}_{\mu}-\alpha\tilde{\mathcal{G}}_{\mu}}(\mathrm{T}_{s},\mathrm{T}_{t})\geq 0$, which is equivalent by Proposition 11 (c2) with $\mathcal{F}-\alpha\mathcal{G}$ convex along $t\mapsto\mu_{t}$. The result for the $\beta$-smoothness follows likewise. ∎

### v2

@@ C.1 Wasserstein differentials
We recall the notion of Wasserstein differentiability introduced in [17, 74, 5]. First, we introduce sub- and super-differentials.
@@
Let $\mathcal{F}:\mathcal{P}_{2}(\mathbb{R}^{d})\to(-\infty,+\infty]$ lower semicontinuous and denote $D(\mathcal{F})=\{\mu\in\mathcal{P}_{2}(\mathbb{R}^{d}),\ \mathcal{F}(\mu)<\infty\}$. Let $\mu\in D(\mathcal{F})$. Then, a map $\xi\in L^{2}(\mu)$ belongs to the subdifferential $\partial^{-}\mathcal{F}(\mu)$ of $\mathcal{F}$ at $\mu$ if for all $\nu\in\mathcal{P}_{2}(\mathbb{R}^{d})$,
@@
$\mathcal{T}_{\mu}\mathcal{P}_{2}(\mathbb{R}^{d})=\overline{\left\{\nabla\psi,\;\psi\in\mathcal{C}_{c}^{\infty}(\mathbb{R}^{d})\right\}}\subset L^{2}(\mu),$ (27)
where the closure is taken in $L^{2}(\mu)$, see Ambrosio et al. [5, Definition 8.4.1]. Lanzetti et al. [74, Proposition 2.5] showed that if $\mathcal{F}$ is Wasserstein differentiable, then there is always a unique gradient living in the tangent space, and we can restrict ourselves without loss of generality to this gradient.
@@
Under regularity assumptions, the Wasserstein gradient of $\mathcal{F}$ can be computed in practice using the first variation $\frac{\delta\mathcal{F}}{\delta\mu}$ [110, Definition 7.12], which is defined, if it exists, as the unique function (up to a constant) such that, for $\chi$ satisfying $\int\mathrm{d}\chi=0$,
@@
Then the Wasserstein gradient can be computed as $\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mu)=\nabla\frac{\delta\mathcal{F}}{\delta\mu}(\mu)$, see e.g. [34, Proposition 5.10].
@@
See Section H.1. ∎
A similar formula can be found in Gangbo and Tudorascu [55, Corollary 3.22], however the space $H$ used there is not $L^{2}(\mu)$ but a lifting $L^{2}(\Omega;\mathbb{R}^{d})$ of measures on random variables. They should not be confused.
@@ C.2 Wasserstein Hessians
A natural object of interest in the context of optimization over the Wasserstein space is the Hessian of the objective $\mathcal{F}$, which we define below, according to the original definitions of [97, Section 3] and [124, Chapter 8]. This notion is usually defined along Wasserstein geodesics.
Definition 8 (Chapter 8 in [124]).
If $\mu$ is absolutely continuous, Wassertein geodesics starting from $\mu$ are curves of the form $\rho_{t}=(\mathrm{Id}+t\nabla\psi)_{\#}\mu$ for $\psi\in\mathcal{C}_{c}^{\infty}(\mathbb{R}^{d})$. Using this fact, one can compute the Wasserstein Hessians of Kullback–Leibler divergence [97], Maximum Mean Discrepancy [7] or Kernel Stein Discrepancy [72] and many other functionals.
However in this work, we are interested in more general curves, which we call acceleration free, i.e. $\mu_{t}=(\mathrm{S}+tv)_{\#}\mu$ with $\mathrm{S},v\in L^{2}(\mu)$. Thus, we define analogously the Hessian along such curves.
@@
We define the Hessian operator $\mathrm{H}\mathcal{F}_{\mu,t}:L^{2}(\mu)\to L^{2}(\mu)$ as the operator satisfying for all $t\in[0,1]$,
where $t\mapsto\mu_{t}=(\mathrm{S}+tv)_{\#}\mu$ for $\mathrm{S},v\in L^{2}(\mu)$.
Note that the Hessian $\mathrm{H}\mathcal{F}_{\mu,t}$ is taken at time $t$ but that the vector field $v\in L^{2}(\mu)$ is in the tangent space at $t=0$, hence the discrepancy with Definition 8 besides the fact that we can have $S\neq\mathrm{Id}$.
Wang and Li [126] derived a general closed form of the Wasserstein Hessian on tangent spaces through the first variation of $\mathcal{F}$. Here, we extend their formula along any curve $\mu_{t}=(\mathrm{S}+tv)_{\#}\mu$ with $\mathrm{S},v\in L^{2}(\mu)$. We first provide a lemma computing the derivative of the Wasserstein gradient.
@@
Let $\mathcal{F}:\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}$ be twice continuously differentiable and assume that $\frac{\delta}{\delta\mu}\nabla\frac{\delta\mathcal{F}}{\delta\mu}(\mu)=\nabla\frac{\delta^{2}\mathcal{F}}{\delta\mu^{2}}(\mu)$ for all $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$. Let $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$ and for all $t\in[0,1]$, $\mu_{t}=(\mathrm{T}_{t})_{\#}\mu$ where $\mathrm{T}_{t}$ is differentiable w.r.t. $t$ with $\frac{\mathrm{d}\mathrm{T}_{t}}{\mathrm{d}t}\in L^{2}(\mu)$. Then,
@@
Under the same assumptions as in Lemma 11, let $\mu_{t}=(\mathrm{T}_{t})_{\#}\mu$ with $\mathrm{T}_{t}=\mathrm{S}+tv$, $\mathrm{S},v\in L^{2}(\mu)$, then $\mathrm{H}\mathcal{F}_{\mu,t}:L^{2}(\mu)\to L^{2}(\mu)$ (as defined in Definition 9) is defined for all $v\in L^{2}(\mu)$, $x\in\mathbb{R}^{d}$ as:
@@
While $\mathrm{H}\mathcal{F}_{\mu,t}$ and $\mathrm{H}\mathcal{F}_{\mu_{t}}$ differ in general, in some simple cases their relation boils down to composition with a pushforward. For instance, if $\mathrm{S}$ is invertible, we can write $\mu_{t}$ as a curve starting from $\mathrm{S}_{\#}\mu$ with a velocity field $v\circ\mathrm{S}^{-1}$, i.e. $\mu_{t}=(\mathrm{Id}+tv\circ\mathrm{S}^{-1})_{\#}\mathrm{S}_{\#}\mu$. Thus, we recover the original definition of the Wasserstein Hessian at $t=0$ as $\mathrm{H}\mathcal{F}_{\mu,0}=\mathrm{H}\mathcal{F}_{\mathrm{S}_{\#}\mu}$. However, in general, this does not need to be the case.color=pink!200,size=] AK: du coup là c’est un peu répétitif avec eq 33 non?color=violet!30,size=] PCA: je suis d’accord mais en même temps dans un case c’est $\mathrm{S}$ qui est inversible, dans l’autre $\mathrm{T}_{t}$ (et ce n’est pas équivalent).
Similarly, if $\mathrm{T}_{t}$ is invertible for all $t$, setting $v_{t}=v\circ\mathrm{T}_{t}^{-1}$, we can write
@@
$=\int\langle\mathrm{H}\mathcal{F}_{\mu,t}[v](x),v(x)\rangle\ \mathrm{d}\mu(x)$
@@
$=\langle\mathrm{H}\mathcal{F}_{\mu_{t}}v_{t},v_{t}\rangle_{L^{2}(\mu_{t})}.$
The last line is obtained by leveraging the closed form in Proposition 12 and that $\mu_{t}=(\mathrm{T}_{t})_{\#}\mu$, as for all $x\in\mathrm{supp}(\mu)$,
$\mathrm{H}\mathcal{F}_{\mu,t}[v](x)$ $=\int\nabla_{y}\nabla_{x}\frac{\delta^{2}\mathcal{F}}{\delta\mu^{2}}(\mu_{t})\big{(}\mathrm{T}_{t}(x),\mathrm{T}_{t}(y)\big{)}v(y)\ \mathrm{d}\mu(y)+\nabla^{2}\frac{\delta\mathcal{F}}{\delta\mu}(\mu_{t})\big{(}\mathrm{T}_{t}(x)\big{)}v(x)$ (34)
$=\int\nabla_{y}\nabla_{x}\frac{\delta^{2}\mathcal{F}}{\delta\mu^{2}}(\mu_{t})(\mathrm{T}_{t}(x),y_{t})v_{t}(y_{t})\ \mathrm{d}\mu_{t}(y)+\nabla^{2}\frac{\delta\mathcal{F}}{\delta\mu}(\mu_{t})\big{(}\mathrm{T}_{t}(x)\big{)}v(x)$
$=\mathrm{H}\mathcal{F}_{\mu_{t}}[v_{t}]\big{(}\mathrm{T}_{t}(x)\big{)},$
and thus $\mathrm{H}\mathcal{F}_{\mu,t}[v]\big{(}\mathrm{T}_{t}^{-1}(x)\big{)}=\mathrm{H}\mathcal{F}_{\mu_{t}}[v_{t}](x)$.
@@
Let $\mathcal{V}(\mu)=\int V\ \mathrm{d}\mu$ with $V$ twice differentiable with bounded Hessian. Then, we have $\frac{\delta\mathcal{V}}{\delta\mu}(\mu)=V$ and $\frac{\delta^{2}\mathcal{V}}{\delta\mu^{2}}(\mu)=0$ (using (29)). Thus, applying Proposition 12, we recover for $\mu_{t}=(\mathrm{T}_{t})_{\#}\mu$,
@@
Let $\mathcal{W}(\mu)=\frac{1}{2}\iint W(x-y)\ \mathrm{d}\mu(x)\mathrm{d}\mu(y)$ with $W$ symmetric, twice differentiable and with bounded Hessian. Then, we have for all $x,y\in\mathbb{R}^{d}$, $\frac{\delta\mathrm{W}}{\delta\mu}(\mu)(x)=(W\star\mu)(x)$ and $\frac{\delta^{2}\mathrm{W}}{\delta\mu^{2}}(\mu)(x,y)=W(x-y)$ (see e.g. [126, Example 7]), and thus applying Proposition 12, for $\mu_{t}=(\mathrm{T}_{t})_{\#}\mu$, the operator is
@@ C.3 Convexity in Wasserstein space
We first recall the definition of $\alpha$-convex functionals [5, Definition 9.1.1].
@@
Let $\mathcal{F}:\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}$ be twice continuously differentiable. Let $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$, $\mathrm{T},\mathrm{S}\in L^{2}(\mu)$, $\mu_{t}=\big{(}\mathrm{T}_{t}\big{)}_{\#}\mu$ for all $t\in[0,1]$ where $\mathrm{T}_{t}=(1-t)\mathrm{S}+t\mathrm{T}$. Furthermore, denote for $t_{1},t_{2}\in[0,1]$, $\tilde{\mu}_{t}^{t_{1}\to t_{2}}=\big{(}(1-t)\mathrm{T}_{t_{1}}+t\mathrm{T}_{t_{2}}\big{)}_{\#}\mu$. Then, the following statement are equivalent:
@@
If we require the convexity and the smoothness to hold along any curve of the form $\mu_{t}=\big{(}(1-t)\mathrm{S}+t\mathrm{T})_{\#}\mu$ for $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$ and $\mathrm{T},\mathrm{S}\in L^{2}(\mu)$, then it coincides with the transport convexity and smoothness recently introduced by Tanaka [117, Definitions 4.1 and 4.5]. As by Proposition 1, $\delta\tilde{\mathcal{F}}_{\mu}(\mathrm{S},\mathrm{T}-\mathrm{S})=\langle\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mathrm{S}_{\#}\mu)\circ\mathrm{S},\mathrm{T}-\mathrm{S}\rangle_{L^{2}(\mu)}$, and thus the convexity of $\mathcal{F}$ on such a curve reads as follows
@@
This type of convexity is actually a particular case of the notion of convexity along acceleration-free curves introduced by Parker [98] (also introduced by Cavagnari et al. [28] under the name of total convexity). The latter requires convexity to hold along any curve of the form $\mu_{t}=\big{(}(1-t)\pi^{1}+t\pi^{2}\big{)}_{\#}\gamma$ with $\gamma\in\Pi(\mu,\nu)$, $\mu,\nu\in\mathcal{P}_{2}(\mathbb{R}^{d})$ and $\pi^{1}(x,y)=x$, $\pi^{2}(x,y)=y$. The transport convexity of Tanaka [117] is thus a particular case for couplings obtained through maps, i.e. $\gamma=(\mathrm{T},\mathrm{S})_{\#}\mu$. Parker [98] notably showed that this notion of convexity is equivalent to the geodesic convexity for Wasserstein differentiable functionals.
We can also define the strict convexity using strict inequalities in Proposition 13-(c1)-(c2)-(c3), but not in (c4) as there are counter examples already for real functions. For instance, $f:\mathbb{R}\to\mathbb{R}$, defined as $f(x)=x^{4}$ for all $x\in\mathbb{R}$, is strictly convex but $f^{\prime\prime}(0)=0$. Thus, for $\mathcal{F}(\mu)=\int f\ \mathrm{d}\mu$, choosing $\mu=\delta_{0}$ and $\mathrm{T}_{0}=\mathrm{Id}$, by Example 1, we have that $\frac{\mathrm{d}^{2}}{\mathrm{d}t^{2}}\mathcal{F}(\mu_{t})\big{|}_{t=0}=f^{\prime\prime}(0)v(0)^{2}=0$. But $\mathcal{F}(\mu_{t})=f\big{(}t\mathrm{T}_{1}(0)\big{)}<tf\big{(}\mathrm{T}_{1}(0)\big{)}$ since $f$ is strictly convex and thus $\mathcal{F}$ is well strictly convex.
Finally, as we defined the relative $\alpha$-convexity and $\beta$-smoothness of $\mathcal{F}$ relative to $\mathcal{G}$ using Bregman divergences in Definition 3, we can show that it is equivalent to $\mathcal{F}-\alpha\mathcal{G}$ and $\beta\mathcal{G}-\mathcal{F}$ being convex.
@@
By Definition 3, $\mathcal{F}$ is $\alpha$-convex relative to $\mathcal{G}$ along $t\mapsto\mu_{t}$ if for all $s,t\in[0,1]$, $\mathrm{d}_{\tilde{\mathcal{F}}_{\mu}}(\mathrm{T}_{s},\mathrm{T}_{t})\geq\alpha\mathrm{d}_{\tilde{\mathcal{G}}_{\mu}}(\mathrm{T}_{s},\mathrm{T}_{t})$. This is equivalent to $\mathrm{d}_{\tilde{\mathcal{F}}_{\mu}-\alpha\tilde{\mathcal{G}}_{\mu}}(\mathrm{T}_{s},\mathrm{T}_{t})\geq 0$, which is equivalent by Proposition 13 (c2) (and using Proposition 1) to $\mathcal{F}-\alpha\mathcal{G}$ convex along $t\mapsto\mu_{t}$. The result for the $\beta$-smoothness follows likewise. ∎